OpenStudy (anonymous):

The social studies teacher is giving a test worth 110 points, and containing 46 questions. There are 2 point and 4 point questions on the test. How many of each type of question are on the test? Show your work.

OpenStudy (mathmate):

Finally, I hope: This is typically a system of equations problem, but I usually solve it by mental reasoning. If all questions are 2-points, we have 2*46=92 points. Since we are 110-92=18 points short, we must have answered 9 4-point questions. So can you figure out the final answers?

OpenStudy (anonymous):

There would be 37 2-point questions and 9 4-point questions

OpenStudy (mathmate):

Did you check the score to make sure it's correct?

OpenStudy (mathmate):

Also, would you be ale to solve a similar problem using the same method?

OpenStudy (anonymous):

No I did not im not sure how to check it

OpenStudy (anonymous):

and yes Im pretty sure I can solve a similar question with this method thank you!

OpenStudy (anonymous):

lol

OpenStudy (mathmate):

Checking: 37*2+9*4=74+36=110 as expected!

OpenStudy (mathmate):

You're welcome! :)

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

Whats so funny @cram

OpenStudy (anonymous):

he told you so it was kinnda like two year olds fighting over a toy

OpenStudy (anonymous):

sorry I am weird that way